# thanks: Jointly supervised this workthanks: Jointly supervised this work

### v1

@@ Acknowledgments
A. S. acknowledges the hospitality of PRL, Ahmedabad, during this work and thanks Amartya Pal for valuable discussions. P. D. acknowledges Department of Space (DoS), India for all the support at PRL, Ahmedabad and, the hospitality at IIT Kanpur during this work. S. K. G. acknowledges financial support from IIT Kanpur via the Initiation Grant (IITK/PHY/2022116). P. D. and S. K. G. also acknowledge Anusandhan National Research Foundation (ANRF) erstwhile Science and Engineering Research Board (SERB), India for financial support through the Start-up Research Grants: SRG/2022/001121 and SRG/2023/000934 respectively.
@@ Appendix A Lindblad formalism to describe dissipation due to the normal lead
To investigate the impact of coupling a ‘lossy’ metallic lead at the junction to an electron reservoir on the transport properties of a TIJJ, we model the system as an open quantum system interacting with its environment, following Ref. [20]. The total Hamiltonian can be decomposed as $H=H_{0}+H_{E}+H_{c}$, where $H_{0}$ is the system Hamiltonian, $H_{E}$ describes the environment, and $H_{c}$ represents the coupling between the system and environment. The dissipative dynamics induced by this coupling term can be described by the Lindblad master equation [26, 27, 46, 47]:
@@
Here, $\rho_{0}$ is the reduced density matrix of the system, ^$\hat{\mathcal{L}}_{m}$ is the Lindblad (quantum jump) operator, and $\eta>0$ is the dissipation rate. Rearranging this equation yields:
@@
The non-Hermitian term $-\frac{i}{2}\sum_{m}\eta\hat{\mathcal{L}}_{m}\hat{\mathcal{L}}^{\dagger}_{m}$ accounts for the ‘loss’ due to coupling with the environment. By neglecting quantum jumps in the second term on the right-hand side of Eq. (12), the system’s dynamics can be effectively described by $H_{\text{eff}}$ [25, 20]. This motivates modeling the non-Hermiticity introduced by the ‘lossy’ lead in the TIJJ by an imaginary potential term $-iV_{2}$ where $V_{2}>0$ is real.
@@ Appendix B Wavefunctions and bound states for NH-TIJJs
The wavefunctions for the two regions $x<0$ ($x_{-}$) and $x>0$ ($x_{+}$) of the NH-TIJJ, within the Andreev approximation and considering the limit of large Fermi energy ($\mu\gg\tilde{\Delta}(\epsilon)$), obtained by solving Eq. (2) have the forms [12],
@@
where, $A$, $B$, $C$ and $D$ are constants, and $k_{F}$ is the Fermi wavevector. The Andreev reflection amplitudes for the electron-like ($+$) and hole-like (--) quasiparticles are,
@@
Inserting Eqs. (14) and (15) in Eq. (3), we get a set of four equations for $A,B,C$ and $D$. For non-trivial solutions, the determinant of the coefficients of this system of equations must vanish leading to the following equation for the ABS energy:
$[a^{+}_{L}-a^{-}_{L}][a^{+}_{R}-a^{-}_{R}]=\mathcal{\zeta}(Z,\theta)[a^{+}_{R}-a^{+}_{L}][a^{-}_{L}-a^{-}_{R}]\;.$ (18)
Substituting the coefficients from Eq. (16) in Eq. (18) we obtain the following equation which can be rearranged to get Eq. (4),
$\left(\frac{\epsilon}{\Delta}+\frac{\epsilon}{\Gamma_{0}}\sqrt{1-\frac{\epsilon^{2}}{\Delta^{2}}}\right)^{2}=1-\zeta(Z,\theta)\sin^{2}\left(\frac{\phi}{2}\right)\;.$ (19)
@@
$c_{1}$ $=$ $-c_{2}=\left[1-\mathcal{\zeta}(Z,\theta)\sin^{2}{\left(\frac{\phi}{2}\right)}\right]^{\frac{1}{2}}\;,$
$c_{3}$ $=$ $\left[1-\mathcal{\zeta}(Z,\theta)\sin^{2}{\left(\frac{\phi}{2}\right)}\right]^{\frac{1}{2}}$ (20)
$+$ $\frac{1}{2}\left[1-\mathcal{\zeta}(Z,\theta)\sin^{2}{\left(\frac{\phi}{2}\right)}\right]^{\frac{3}{2}}.$
Figure 5: ABS energy for purely non-Hermitian TIJJ ($Z_{1}=0$, $\theta=\pi/6$): The complex ABS energy plotted for two different values of $Z_{2}$ in (a) and (b) taking $\gamma=0.3$, and for two different values of $\gamma$ in (c) and (d) keeping $Z_{2}=0.1$.
For the purely non-Hermitian case ($Z_{1}=0$, $Z_{2}\neq 0$), the variation of the ABS energies in this NH-TIJJ for different values of $Z_{2}$ and $\gamma$ is illustrated in Fig. 5. Fig. 5(a) and (b) show that the imaginary component of the ABS energy increases with increasing $Z_{2}$. Hence, as the strength of non-Hermiticity grows, the lifetime of the ABS quasiparticles in the junction decreases, leading to greater decoherence in charge transport. Furthermore, the length of the zero-energy line increases with $Z_{2}$, indicating that non-Hermiticity directly facilitates the formation of these ZES. The variation of the ABS energy with $\gamma$ in this regime is shown in Fig. 5(c) and (d). Increasing $\gamma$ raises the maximum energy of the bound states, as enhanced tunneling allows more quasiparticles to traverse from the superconducting regions into the junction, resulting in a larger number of bound states. However, the zero-energy line remains unaffected by changes in $\gamma$.

### v2

@@ Acknowledgments
A. S. acknowledges the hospitality of PRL, Ahmedabad, during this work and thanks Amartya Pal and D. Michel Pino for valuable discussions. P. D. acknowledges Department of Space (DoS), India for all the support at PRL, Ahmedabad and, the hospitality at IIT Kanpur during this work. S. K. G. acknowledges financial support from IIT Kanpur via the Initiation Grant (IITK/PHY/2022116). P. D. and S. K. G. also acknowledge Anusandhan National Research Foundation (ANRF) erstwhile Science and Engineering Research Board (SERB), India for financial support through the Start-up Research Grants: SRG/2022/001121 and SRG/2023/000934 respectively.
@@ Appendix A Lindblad formalism to describe dissipation due to the normal lead
To investigate the impact of coupling a ‘lossy’ metallic lead at the junction to an electron reservoir on the transport properties of a TIJJ, we model the system as an open quantum system interacting with its environment, following Ref. [Li2024]. The total Hamiltonian can be decomposed as $H=H_{0}+H_{E}+H_{c}$, where $H_{0}$ is the system Hamiltonian, $H_{E}$ describes the environment, and $H_{c}$ represents the coupling between the system and environment. The dissipative dynamics induced by this coupling term can be described by the Lindblad master equation [breuer2002, daley2014, roccati2022non, Manzano2020]:
@@
Here, $\rho_{0}$ is the reduced density matrix of the system, $\hat{\mathcal{L}}_{m}$ is the Lindblad (quantum jump) operator, and $\eta>0$ is the dissipation rate. Rearranging this equation yields:
@@
The non-Hermitian term $\frac{i}{2}\sum_{m}\eta\hat{\mathcal{L}}_{m}\hat{\mathcal{L}}^{\dagger}_{m}$ accounts for the ‘loss’ due to coupling with the environment. By neglecting quantum jumps in the second term on the right-hand side of Eq. (13), the system’s dynamics can be effectively described by $H_{\text{eff}}$ [Harrington2022, Li2024]. This motivates modeling the non-Hermiticity introduced by the ‘lossy’ lead in the TIJJ by an imaginary potential term $iV_{2}$ where $V_{2}>0$ is real.
Figure 6: Characteristics of the NH-TIJJ in the strong-proximity limit ($\Gamma_{0}=\Delta$): (a) The real (solid lines) and imaginary (dashed lines) parts of the Andreev spectrum of the NH-TIJJ as a function of the phase difference $\phi$ plotted for different angles of incidence $\theta$. (b) Variation of the imaginary part of the spectrum as a function of $\phi$ and $Z_{2}$ for the incident angle $\theta=\pi/3$, the boundary of the dark purple region demonstrates the asymmetric gap-exit points in the spectrum. (c) The angle-resolved current $I_{\theta}(\phi)$ plotted as a function of $\phi$ for different angles of incidence. (d) The total supercurrent $I_{\mathrm{total}}(\phi)$ as a function of $\phi$ for different barrier parameters. The current is computed in units of $e\Delta/\hbar$ and we have chosen the real barrier parameter $Z_{1}=0.4$ and the imaginary barrier parameter $Z_{2}=0.2$
@@ Appendix B Wavefunctions and bound states for NH-TIJJs
The wavefunctions for the two regions $x<0$ ($x_{-}$) and $x>0$ ($x_{+}$) of the NH-TIJJ, within the Andreev approximation and considering the limit of large Fermi energy ($\mu\gg\tilde{\Delta}(\epsilon)$), obtained by solving Eq. (2) have the forms [Tkachov2013],
@@
where, $A$, $B$, $C$ and $D$ are constants, and $k_{F}$ is the Fermi wavevector. The Andreev reflection amplitudes for the electron-like (++) and hole-like (-) quasiparticles are,
@@
Inserting Eqs. (15) and (16) in Eq. (3), we get a set of four equations for $A,B,C$ and $D$. For non-trivial solutions, the determinant of the coefficients of this system of equations must vanish, leading to the following equation for the ABS energy:
$T_{1}(a^{+}_{L}a^{-}_{L}+a^{+}_{R}a^{-}_{R})-T_{2}(a^{+}_{L}a^{+}_{R}+a^{-}_{L}a^{-}_{R})+$
$T_{3}(a^{-}_{L}a^{+}_{R}+a^{+}_{L}a^{-}_{R})-2iZ_{1}Z_{2}\cos^{2}{(\theta)}(a^{+}_{L}a^{-}_{L}-a^{+}_{R}a^{-}_{R})$
$+2Z_{2}\cos{(\theta)}(a^{+}_{L}a^{+}_{R}-a^{-}_{L}a^{-}_{R})=0\;.$ (19)
where $T_{1}=(1+Z^{2}_{1}-Z^{2}_{2})\cos^{2}{(\theta)}$, $T_{2}=Z^{2}_{1}+Z^{2}_{2}+\cos^{2}{(\theta)}$ and $T_{3}=\sin^{2}{(\theta)}(Z^{2}_{1}+Z^{2}_{2})$. Substituting the coefficients from Eq. (17) in Eq. (B) we arrive at Eq. (4). Now from Eq. (5), denoting the R.H.S as $\mathcal{S}(\theta,\phi)$ and simplifying the L.H.S we obtain the following equation which can be rearranged to get Eq. (6),
$\left(\frac{\epsilon}{\Delta}+\frac{\epsilon}{\Gamma_{0}}\sqrt{1-\frac{\epsilon^{2}}{\Delta^{2}}}\right)^{2}=\mathcal{S}(\theta,\phi)\;.$ (20)
@@
$c_{1}$ $=$ $-c_{2}=\mathcal{S}^{\frac{1}{2}}(\theta,\phi),$
$c_{3}$ $=$ $\mathcal{S}^{\frac{1}{2}}(\theta,\phi)+\frac{1}{2}\mathcal{S}^{\frac{3}{2}}(\theta,\phi).$ (21)
For the NH-TIJJ in the strong-proximity limit ($\Gamma_{0}=\Delta$), the Andreev spectrum is shown in Fig. 6(a) and (b) and bears similarities in qualitative behaviors with that of the tunneling limit (see Fig. 2). This includes features like the spectrum bifurcation and asymmetry of the GEPs, note that for the strong-proximity limit the bifurcation happens around the parent superconducting gap $\Delta$ with one branch going above $\Delta$ and another below it, whereas in the tunneling limit the bifurcation occurs around the induced gap which is much smaller than $\Delta$. Similarly, the angle-resolved current (Fig. 6(c)) and the total current CPR (Fig. 6(d)) exhibit the anomalous behavior stemming from the asymmetry of the spectrum.